Grade 12Mechanics

A block A slides down a smooth inclined plane when released from the top in time t. Another block falls freely from the same point and strikes the ground in time t/2. The value of angle teta is

mrli
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear Student
 

If angle of inclination is θ, then the acceleration with which the block slides on the incline is g sin \theta.

Given that the block is taking a time t/2 and when it is left freely falling from the same height, then the block falls with an acceleration g, at this time it is taking a time t. therefore

S = u* t + (1/2)*g *t^2
h = 0 – (1/2)*g sin\theta * t^2
 
t^2 = 2 h/ g sin\theta
and 
h sin\theta = (1/2)*g *(t/2)^2
t^2 = 8 h sin\theta / g
 
Now
8 h sin\theta / g = 2 h/ g sin\theta
sin2\theta = ¼
sin\theta = ½
\theta = 30 degree
 
Regards
Arun (askIITians forum expert)